# Deep FlexQP: Accelerated Nonlinear Programming via Deep Unfolding

### v1

@@ 6 Conclusion
Next, let $\hat{x}$ solve Equation 15. If $x^{*}\neq\hat{x}$ solves Equation 3, then we have that
$\phi(\hat{x};\mu_{I},\mu_{E})$ $=\frac{1}{2}\hat{x}^{\top}P\hat{x}+q^{\top}\hat{x}+\mu_{I}\left\lVert(G\hat{x}-h)_{+}\right\rVert_{1}+\mu_{E}\left\lVert A\hat{x}-b\right\rVert_{1}$ (17)
$\leq\frac{1}{2}{x^{*}}^{\top}Px^{*}+q^{\top}x^{*}+\mu_{I}\left\lVert(Gx^{*}-h)_{+}\right\rVert_{1}+\mu_{E}\left\lVert Ax^{*}-b\right\rVert_{1}$ (18)
$=\frac{1}{2}{x^{*}}^{\top}Px^{*}+q^{\top}x^{*},$ (19)
which follows by the optimality of $\hat{x}$. Now, assume that $\hat{x}$ is not feasible for Equation 3. Then
$\frac{1}{2}\hat{x}^{\top}P\hat{x}+q^{\top}\hat{x}$ $\geq\frac{1}{2}{x^{*}}^{\top}Px^{*}+q^{\top}x^{*}+(Px^{*}+q)^{\top}(\hat{x}-x^{*})$ (20)
$=\frac{1}{2}{x^{*}}^{\top}Px^{*}+q^{\top}x^{*}-\sum_{i=1}^{m}y_{I,i}^{*}g_{i}^{\top}(\hat{x}-x^{*})-\sum_{i=1}^{p}y_{E,i}^{*}a_{i}^{\top}(\hat{x}-x^{*})$ (21)
$=\begin{aligned} \frac{1}{2}{x^{*}}^{\top}Px^{*}+q^{\top}x^{*}&-\sum_{i=1}^{m}y_{I,i}^{*}(g_{i}^{\top}\hat{x}-h_{i}-(g_{i}^{\top}x^{*}-h_{i}))\\ &\hskip 20.00003pt-\sum_{i=1}^{p}y_{E,i}^{*}(a_{i}^{\top}\hat{x}-b_{i}-(a_{i}^{\top}x^{*}-b_{i}))\end{aligned}$ (22)
$=\frac{1}{2}{x^{*}}^{\top}Px^{*}+q^{\top}x^{*}-\sum_{i=1}^{m}y_{I,i}^{*}(g_{i}^{\top}\hat{x}-h_{i})-\sum_{i=1}^{p}y_{E,i}^{*}(a_{i}^{\top}\hat{x}-b_{i})$ (23)
$\geq\frac{1}{2}{x^{*}}^{\top}Px^{*}+q^{\top}x^{*}-\mu_{I}\sum_{i=1}^{m}g_{i}^{\top}\hat{x}-h_{i}-\mu_{E}\sum_{i=1}^{p}a_{i}^{\top}\hat{x}-b_{i}$ (24)
$\geq\frac{1}{2}{x^{*}}^{\top}Px^{*}+q^{\top}x^{*}-\mu_{I}\sum_{i=1}^{m}(g_{i}^{\top}\hat{x}-h_{i})_{+}-\mu_{E}\sum_{i=1}^{p}|a_{i}^{\top}\hat{x}-b_{i}|,$ (25)
@@
$\frac{1}{2}\hat{x}^{\top}P\hat{x}+q^{\top}\hat{x}+\mu_{I}\left\lVert(G\hat{x}-h)_{+}\right\rVert_{1}+\mu_{E}\left\lVert Ax-b\right\rVert_{1}\geq\frac{1}{2}{x^{*}}^{\top}Px^{*}+q^{\top}x^{*},$ (26)
but either $\hat{x}=x^{*}$ or this contradicts the fact that $\hat{x}$ minimized $\phi(\cdot;\mu_{I},\mu_{E})$ in Equation 19. Thus, $\hat{x}$ must be feasible for Equation 3. Therefore, by Equation 19 we have that
@@
so $\hat{x}$ minimizes the quadratic objective and thus solves Equation 3, completing the proof.
Appendix B FlexQP — First Block ADMM Update
The most computationally demanding step of FlexQP is the first block update Equation 9a, which is an equality-constrained QP:
@@ Appendix B FlexQP — First Block ADMM Update
where $\tilde{\nu}_{I}\in{\mathbb{R}}^{m}$ and $\tilde{\nu}_{E}\in{\mathbb{R}}^{p}$ are the Lagrange multipliers for the equality constraints Equation 28b. Solving this QP as-is would be expensive since it requires solving a linear system of size $n+3m+2p$. However, we can eliminate $\tilde{s}$, $\tilde{z}_{I}$, and $\tilde{z}_{E}$ using Equations 29b, 29c and 29d above, so the linear system simplifies to
@@
The coefficient matrix in the linear system Equation 30 is always full rank due to the positive parameters $\sigma_{x}$, $\sigma_{s}$, $\rho_{I}$, and $\rho_{E}$ introduced through the ADMM splitting. This linear system can be solved using a direct method such as an $LDL^{\top}$ factorization requiring $O((n+m+p)^{3})$ time, the same as OSQP using the direct method. On the other hand, for large-scale QPs, i.e., when $n+m+p$ is very large, factoring this matrix can be prohibitively expensive. In this case, we can use an indirect method to solve the reduced system
$\begin{aligned} (P+\sigma_{x}I+\bar{G}^{\top}G+\bar{A}^{\top}A)\tilde{x}={}&\sigma_{x}x^{k}-q+\bar{G}^{\top}(h-s^{k}+\sigma_{s}^{-1}w_{s}^{k}+z_{I}^{k}-\rho_{I}^{-1}y_{I}^{k})\\ &+\bar{A}^{\top}(b+z_{E}^{k}-\rho_{E}^{-1}y_{E}^{k}),\end{aligned}$ (32)
where $\bar{G}=(\sigma_{s}^{-1}+\rho_{I}^{-1})^{-1}G$ and $\bar{A}=\rho_{E}A$. This can be obtained by eliminating $\tilde{\nu}_{I}$ and $\tilde{\nu}_{E}$ from the linear system Equation 30. These variables are recoverable using
@@
The coefficient matrix in Equation 32 is always positive definite, so the linear system can be solved using an iterative algorithm such as the conjugate gradient (CG) method. The linear system is of size $n$, matching the complexity of OSQP using the indirect method. In this work, we consider a supervised learning setting where we will need to compute derivatives of the solution $\tilde{x}$ with respect to the parameters $\sigma_{x}$, $\rho_{I}$, etc. While each iteration of the CG method is very fast, it can require many iterations to converge to a low-error solution. It would be very inefficient to backpropagate through all these iterations of the CG method, the main issue being the high memory cost since the entire compute graph needs to be stored and then differentiated through during the backward pass. We instead adopt the approach from saravanos2025deep using differentiable optimization in order to compute these derivatives in a more efficient manner. In practice, this means we can compute the derivatives by solving a new linear system with the same coefficient matrix but different right-hand side during the backward pass.
Appendix C FlexQP Algorithm
Inputs: Initialization $x^{0},s^{0},z_{I}^{0},z_{E}^{0},w_{s}^{0},y_{I}^{0},y_{E}^{0}$
   and penalty parameters $\mu_{I},\mu_{E},\sigma_{x},\sigma_{s},\rho_{I},\rho_{E}>0$
@@ Appendix C FlexQP Algorithm
 $\tilde{x}^{k+1},\tilde{\nu}_{I}^{k+1},\tilde{\nu}_{E}^{k+1}\leftarrow$ Solve the linear system Equation 30
@@
 $w_{s}^{k+1}=w_{s}^{k}+\sigma_{s}(\tilde{s}^{k+1}-s^{k+1})$
 $y_{I}^{k+1}=y_{I}^{k}+\rho_{I}(\tilde{z}_{I}^{k+1}-z_{I}^{k+1})$
 $y_{E}^{k+1}=y_{E}^{k}+\rho_{E}(\tilde{z}_{E}^{k+1}-z_{E}^{k+1})$
@@
Appendix D Proof of Theorem 3.2
To show that a saddle point of Equation 10 exists, it suffices to show that the relative interior of the domains of $f$ and $g$ are non-empty. This is a form of constraint qualification guaranteeing strong duality (rockafellar1970convex, Theorem 16.4). The relative interior of $f$ is simply the feasible set $\{(x,s,z_{I},z_{E}):z_{I}=Gx+s-h,z_{E}=Ax-b\}$, which is non-empty (pick any $x,s$ and then let $z_{I}=Gx+s-h$ and $z_{E}=Ax-b$). Meanwhile, the relative interior of $g$ is given by the set $\{(x,s,z_{I},z_{E}):s>0\}$ due to the constraint $s\geq 0$. Combining, this shows that $\operatorname{relint}(\operatorname{dom}f)\cap\operatorname{relint}(\operatorname{dom}g)\neq\emptyset$, so strong duality holds and a saddle point exists by rockafellar1970convex.
Furthermore, as the objective $f$ and $g$ are closed, proper, and convex, by boyd2011distributed, Algorithm 1 converges. Namely, we have that the ADMM primal residuals $\tilde{\bm{\zeta}}^{k}\to 0$ and ADMM dual residuals $\bar{\bm{\zeta}}^{k}\to 0$ as $k\to\infty$, where
$\tilde{\bm{\zeta}}^{k}=\begin{bmatrix}\tilde{\zeta}_{x}^{k}\\ \tilde{\zeta}_{s}^{k}\\ \tilde{\zeta}_{I}^{k}\\ \tilde{\zeta}_{E}^{k}\end{bmatrix}=\begin{bmatrix}\tilde{x}^{k}-x^{k}\\ \tilde{s}^{k}-s^{k}\\ \tilde{z}_{I}^{k}-z_{I}^{k}\\ \tilde{z}_{E}^{k}-z_{E}^{k}\end{bmatrix},\quad\bar{\bm{\zeta}}=\begin{bmatrix}\bar{\zeta}_{x}^{k}\\ \bar{\zeta}_{s}^{k}\\ \bar{\zeta}_{I}^{k}\\ \bar{\zeta}_{E}^{k}\end{bmatrix}=\begin{bmatrix}x^{k-1}-x^{k}\\ s^{k-1}-s^{k}\\ z_{I}^{k-1}-z_{I}^{k}\\ z_{E}^{k-1}-z_{E}^{k}\end{bmatrix}$. (34)
@@
Appendix E Proof of Theorem 3.3
The proof follows from the definition of the soft thresholding operator. First, we consider the $z_{I}$ and $z_{E}$ updates from Equations 9d and 9e as well as the dual variable updates for $y_{I}$ and $y_{E}$ from Equations 9h and 9i. Assume w.l.o.g. that $\alpha=1$. These updates have the general form:
$z^{k+1}=S_{\mu/\rho}\left(\tilde{z}^{k+1}+y^{k}/\rho\right),$ (35)
$y^{k+1}=y^{k}+\rho(\tilde{z}^{k+1}-z^{k+1}).$ (36)
@@
    Positive constraint violation: If $\tilde{z}^{k+1}+y^{k}/\rho>\mu/\rho$, then $z^{k+1}=\tilde{z}^{k+1}+y^{k}/\rho-\mu/\rho$. Substituting into Equation 36 yields $y^{k+1}=\mu$.
@@
    No constraint violation: If $|\tilde{z}^{k+1}+y^{k}/\rho|\leq\mu/\rho$, then $z^{k+1}=0$. This further implies $\rho|\tilde{z}^{k+1}+y^{k}/\rho|\leq\mu$ and by Equation 36 this implies $|y^{k+1}|\leq\mu$.
@@
    Negative constraint violation: If $\tilde{z}^{k+1}+y^{k}/\rho<\mu/\rho$, then $z^{k+1}=\tilde{z}^{k+1}+y^{k}/\rho+\mu/\rho$. Substituting into Equation 36 yields $y^{k+1}=-\mu$.
@@
Appendix F Deep FlexQP Policy Parameterization
The residuals for Equation 7 are given by
@@ Appendix F Deep FlexQP Policy Parameterization
The ADMM residuals for Equation 8 are defined in Equation 34. We ignore the residuals corresponding to $x$ since it is unconstrained in the second ADMM block (so the primal residual is not very meaningful) and we have already captured the optimality through Equation 37.
@@
$\alpha=\pi_{\alpha}(\lVert\zeta_{\text{dual}}\rVert,\lVert\zeta_{I}\rVert,\lVert\zeta_{E}\rVert,\lVert\bar{\zeta}_{s}\rVert,\lVert\bar{\zeta}_{I}\rVert,\lVert\bar{\zeta}_{E}\rVert,\lVert\tilde{\zeta}_{s}\rVert,\lVert\tilde{\zeta}_{I}\rVert,\lVert\tilde{\zeta}_{s}\rVert),$ (40)
@@
We consider two policy parameterizations: multilayer perceptrons (MLPs) and LSTMs (see Appendix M for experimental comparison). Following saravanos2025deep, MLP policies are small networks with two hidden layers of sizes [32, 32]. LSTM policies use a hidden size of 32 followed by an MLP with hidden layers [32, 32] for prediction. We use sigmoid for all activation functions, which we find are much more stable than ReLU activations, most likely due to the autoregressive nature of deep unfolding. Computationally, we log-scale any small positive inputs like the infinity norms of the residuals. Following ichnowski2021accelerating, we also predict log-transformed values $\log\mu_{I}$, $\log\rho_{E}$, etc. and then apply an exponential function so that it is easier to predict parameters across a wide scale of values. We then clamp the parameters (besides $\alpha$) to the range [1e6,1e6][\$110-6\$,\$1106\$]; $\alpha\in(0,2)$ is enforced using a scaled sigmoid function.
Appendix G Further Details on QP Problem Classes
A summary of the problem sizes and training parameters of the different classes is presented in Table 1. For the small- to medium-scale QPs, we train all models for 500 epochs and evaluate using 1000 test samples. More training samples are used for random QPs following the setup by saravanos2025deep since the shared structure between these QPs is less clear and therefore harder to learn. As described in the main text, for the large-scale problems, we fine-tune the models trained on the smaller scale problems on 100 large-scale problems for 5 epochs, and test on 100 new problems.

### v2

@@ Appendix A Extended Related Work
Next, let $\hat{x}$ solve Eq. 15. If $x^{*}\neq\hat{x}$ solves Eq. 3, then we have that
$\phi(\hat{x};\mu_{I},\mu_{E})$ $=\frac{1}{2}\hat{x}^{\top}P\hat{x}+q^{\top}\hat{x}+\mu_{I}\left\lVert(G\hat{x}-h)_{+}\right\rVert_{1}+\mu_{E}\left\lVert A\hat{x}-b\right\rVert_{1}$ (17a)
$\leq\frac{1}{2}{x^{*}}^{\top}Px^{*}+q^{\top}x^{*}+\mu_{I}\left\lVert(Gx^{*}-h)_{+}\right\rVert_{1}+\mu_{E}\left\lVert Ax^{*}-b\right\rVert_{1}$ (17b)
$=\frac{1}{2}{x^{*}}^{\top}Px^{*}+q^{\top}x^{*},$ (17c)
which follows by the optimality of $\hat{x}$. Now, assume that $\hat{x}$ is not feasible for Eq. 3. Then
$\frac{1}{2}\hat{x}^{\top}P\hat{x}+q^{\top}\hat{x}$ $\geq\frac{1}{2}{x^{*}}^{\top}Px^{*}+q^{\top}x^{*}+(Px^{*}+q)^{\top}(\hat{x}-x^{*})$ (18a)
$=\frac{1}{2}{x^{*}}^{\top}Px^{*}+q^{\top}x^{*}-\sum_{i=1}^{m}y_{I,i}^{*}g_{i}^{\top}(\hat{x}-x^{*})-\sum_{i=1}^{p}y_{E,i}^{*}a_{i}^{\top}(\hat{x}-x^{*})$ (18b)
$=\begin{aligned} \frac{1}{2}{x^{*}}^{\top}Px^{*}+q^{\top}x^{*}&-\sum_{i=1}^{m}y_{I,i}^{*}(g_{i}^{\top}\hat{x}-h_{i}-(g_{i}^{\top}x^{*}-h_{i}))\\ &\hskip 20.00003pt-\sum_{i=1}^{p}y_{E,i}^{*}(a_{i}^{\top}\hat{x}-b_{i}-(a_{i}^{\top}x^{*}-b_{i}))\end{aligned}$ (18c)
$=\frac{1}{2}{x^{*}}^{\top}Px^{*}+q^{\top}x^{*}-\sum_{i=1}^{m}y_{I,i}^{*}(g_{i}^{\top}\hat{x}-h_{i})-\sum_{i=1}^{p}y_{E,i}^{*}(a_{i}^{\top}\hat{x}-b_{i})$ (18d)
$\geq\frac{1}{2}{x^{*}}^{\top}Px^{*}+q^{\top}x^{*}-\mu_{I}\sum_{i=1}^{m}g_{i}^{\top}\hat{x}-h_{i}-\mu_{E}\sum_{i=1}^{p}a_{i}^{\top}\hat{x}-b_{i}$ (18e)
$\geq\frac{1}{2}{x^{*}}^{\top}Px^{*}+q^{\top}x^{*}-\mu_{I}\sum_{i=1}^{m}(g_{i}^{\top}\hat{x}-h_{i})_{+}-\mu_{E}\sum_{i=1}^{p}|a_{i}^{\top}\hat{x}-b_{i}|,$ (18f)
@@
$\frac{1}{2}\hat{x}^{\top}P\hat{x}+q^{\top}\hat{x}+\mu_{I}\left\lVert(G\hat{x}-h)_{+}\right\rVert_{1}+\mu_{E}\left\lVert A\hat{x}-b\right\rVert_{1}\geq\frac{1}{2}{x^{*}}^{\top}Px^{*}+q^{\top}x^{*},$ (19)
but either $\hat{x}=x^{*}$ or this contradicts the fact that $\hat{x}$ minimized $\phi(\cdot;\mu_{I},\mu_{E})$ in Eq. 17c. Thus, $\hat{x}$ must be feasible for Eq. 3. Therefore, by Eq. 17c we have that
@@
so $\hat{x}$ minimizes the quadratic objective and thus solves Eq. 3, completing the proof.
Appendix C FlexQP — First Block ADMM Update
The most computationally demanding step of FlexQP is the first block update Eq. 9, which is an equality-constrained QP:
@@ Appendix C FlexQP — First Block ADMM Update
where $\tilde{\nu}_{I}\in{\mathbb{R}}^{m}$ and $\tilde{\nu}_{E}\in{\mathbb{R}}^{p}$ are the Lagrange multipliers for the equality constraints Eq. 21b. Solving this QP as-is would be expensive since it requires solving a linear system of size $n+3m+2p$. However, we can eliminate $\tilde{s}$, $\tilde{z}_{I}$, and $\tilde{z}_{E}$ using Eqs. 22b, 22c and 22d above, so the linear system simplifies to
@@
The coefficient matrix in the linear system Eq. 23 is always full rank due to the positive parameters $\sigma_{x}$, $\sigma_{s}$, $\rho_{I}$, and $\rho_{E}$ introduced through the ADMM splitting. This linear system can be solved using a direct method such as an $LDL^{\top}$ factorization requiring $O((n+m+p)^{3})$ time, the same as OSQP using the direct method. On the other hand, for large-scale QPs, i.e., when $n+m+p$ is very large, factoring this matrix can be prohibitively expensive. In this case, we can use an indirect method to solve the reduced system
$(P+\sigma_{x}I+\bar{G}^{\top}G+\bar{A}^{\top}A)\tilde{x}$={} $\sigma_{x}x^{k}-q+\bar{G}^{\top}(h-s^{k}+\sigma_{s}^{-1}w_{s}^{k}+z_{I}^{k}-\rho_{I}^{-1}y_{I}^{k})$ (25)
$+\bar{A}^{\top}(b+z_{E}^{k}-\rho_{E}^{-1}y_{E}^{k}),$
where $\bar{G}=(\sigma_{s}^{-1}+\rho_{I}^{-1})^{-1}G$ and $\bar{A}=\rho_{E}A$. This can be obtained by eliminating $\tilde{\nu}_{I}$ and $\tilde{\nu}_{E}$ from the linear system Eq. 23. These variables are recoverable using
@@
The coefficient matrix in Eq. 25 is always positive definite, so the linear system can be solved using an iterative algorithm such as the conjugate gradient (CG) method. The linear system is of size $n$, matching the complexity of OSQP using the indirect method. In this work, we consider a supervised learning setting where we will need to compute derivatives of the solution $\tilde{x}$ with respect to the parameters $\sigma_{x}$, $\rho_{I}$, etc. While each iteration of the CG method is very fast, it can require many iterations to converge to a low-error solution. It would be very inefficient to backpropagate through all these iterations of the CG method, the main issue being the high memory cost since the entire compute graph needs to be stored and then differentiated through during the backward pass. We instead adopt the approach from Saravanos et al. (2025, Theorem 2) using differentiable optimization in order to compute these derivatives in a more efficient manner. In practice, this means we can compute the derivatives by solving a new linear system with the same coefficient matrix but different right-hand side during the backward pass.
Appendix D FlexQP Algorithm
Input : Initialization $x^{0},s^{0},z_{I}^{0},z_{E}^{0},w_{s}^{0},y_{I}^{0},y_{E}^{0}$ and parameters $\mu_{I},\mu_{E},\sigma_{x},\sigma_{s},\rho_{I},\rho_{E}>0$
@@ Appendix D FlexQP Algorithm
 $\tilde{x}^{k+1},\tilde{\nu}_{I}^{k+1},\tilde{\nu}_{E}^{k+1}\leftarrow$ Solve the linear system Eq. 23
@@
 $w_{s}^{k+1}=w_{s}^{k}+\sigma_{s}(\alpha\tilde{s}^{k+1}+(1-\alpha)s^{k}-s^{k+1})$
 $y_{I}^{k+1}=y_{I}^{k}+\rho_{I}(\alpha\tilde{z}_{I}^{k+1}+(1-\alpha)z_{I}^{k}-z_{I}^{k+1})$
 $y_{E}^{k+1}=y_{E}^{k}+\rho_{E}(\alpha\tilde{z}_{E}^{k+1}+(1-\alpha)z_{E}^{k}-z_{E}^{k+1})$
@@
Appendix E Proof of Theorem 3.2
To show that a saddle point of Eq. 10 exists, it suffices to show that the relative interior of the domains of $f$ and $g$ are non-empty. This is a form of constraint qualification guaranteeing strong duality (Rockafellar, 1970, Theorem 16.4). The relative interior of $f$ is simply the feasible set $\{(x,s,z_{I},z_{E}):z_{I}=Gx+s-h,z_{E}=Ax-b\}$, which is non-empty (pick any $x,s$ and then let $z_{I}=Gx+s-h$ and $z_{E}=Ax-b$). Meanwhile, the relative interior of $g$ is given by the set $\{(x,s,z_{I},z_{E}):s>0\}$ due to the constraint $s\geq 0$. Combining, this shows that $\operatorname{relint}(\operatorname{dom}f)\cap\operatorname{relint}(\operatorname{dom}g)\neq\emptyset$, so strong duality holds and a saddle point exists by Rockafellar (1970, Theorem 37.3).
Furthermore, as the objective $f$ and $g$ are closed, proper, and convex, by Boyd et al. (2011, §3.2), Algorithm 1 converges. Namely, we have that the ADMM primal residuals $\tilde{\bm{\zeta}}^{k}\to 0$ and ADMM dual residuals $\bar{\bm{\zeta}}^{k}\to 0$ as $k\to\infty$, where
$\tilde{\bm{\zeta}}^{k}=\begin{bmatrix}\tilde{\zeta}_{x}^{k}\\ \tilde{\zeta}_{s}^{k}\\ \tilde{\zeta}_{I}^{k}\\ \tilde{\zeta}_{E}^{k}\end{bmatrix}=\begin{bmatrix}\tilde{x}^{k}-x^{k}\\ \tilde{s}^{k}-s^{k}\\ \tilde{z}_{I}^{k}-z_{I}^{k}\\ \tilde{z}_{E}^{k}-z_{E}^{k}\end{bmatrix},\quad\bar{\bm{\zeta}}^{k}=\begin{bmatrix}\bar{\zeta}_{x}^{k}\\ \bar{\zeta}_{s}^{k}\\ \bar{\zeta}_{I}^{k}\\ \bar{\zeta}_{E}^{k}\end{bmatrix}=\begin{bmatrix}x^{k-1}-x^{k}\\ s^{k-1}-s^{k}\\ z_{I}^{k-1}-z_{I}^{k}\\ z_{E}^{k-1}-z_{E}^{k}\end{bmatrix}$. (27)
@@
Appendix F Proof of Theorem 3.3
The proof follows from the definition of the soft thresholding operator. First, we consider the $z_{I}$ and $z_{E}$ updates as well as the dual variable updates for $y_{I}$ and $y_{E}$ from Eq. 9. Assume w.l.o.g. that $\alpha=1$. These updates have the general form:
$z^{k+1}=S_{\mu/\rho}\left(\tilde{z}^{k+1}+y^{k}/\rho\right),$ (28a)
$y^{k+1}=y^{k}+\rho(\tilde{z}^{k+1}-z^{k+1}).$ (28b)
@@
    Positive constraint violation: If $\tilde{z}^{k+1}+y^{k}/\rho>\mu/\rho$, then $z^{k+1}=\tilde{z}^{k+1}+y^{k}/\rho-\mu/\rho$. Substituting into Eq. 28b yields $y^{k+1}=\mu$.
@@
    No constraint violation: If $|\tilde{z}^{k+1}+y^{k}/\rho|\leq\mu/\rho$, then $z^{k+1}=0$. This further implies $\rho|\tilde{z}^{k+1}+y^{k}/\rho|\leq\mu$ and by Eq. 28b this implies $|y^{k+1}|\leq\mu$.
@@
    Negative constraint violation: If $\tilde{z}^{k+1}+y^{k}/\rho<-\mu/\rho$, then $z^{k+1}=\tilde{z}^{k+1}+y^{k}/\rho+\mu/\rho$. Substituting into Eq. 28b yields $y^{k+1}=-\mu$.
@@
Appendix G Deep FlexQP Policy Parameterization
The residuals for Eq. 7 are given by
@@ Appendix G Deep FlexQP Policy Parameterization
The ADMM residuals for Eq. 8 are defined in Eq. 27. We ignore the residuals corresponding to $x$ since it is unconstrained in the second ADMM block (so the primal residual is not very meaningful) and we have already captured the optimality through Eq. 29.
@@
$\alpha=\pi_{\alpha}(\lVert\zeta_{\text{dual}}\rVert,\lVert\zeta_{I}\rVert,\lVert\zeta_{E}\rVert,\lVert\bar{\zeta}_{s}\rVert,\lVert\bar{\zeta}_{I}\rVert,\lVert\bar{\zeta}_{E}\rVert,\lVert\tilde{\zeta}_{s}\rVert,\lVert\tilde{\zeta}_{I}\rVert,\lVert\tilde{\zeta}_{E}\rVert),$ (32)
@@
We consider two policy parameterizations: multilayer perceptrons (MLPs) and LSTMs (see Appendix N for experimental comparison). Following Saravanos et al. (2025), MLP policies are small networks with two hidden layers of sizes [32, 32]. LSTM policies use a hidden size of 32 followed by an MLP with hidden layers [32, 32] for prediction. We use sigmoid for all activation functions, which we find are much more stable than ReLU activations, most likely due to the autoregressive nature of deep unfolding. Computationally, we log-scale any small positive inputs like the infinity norms of the residuals. Following Ichnowski et al. (2021), we also predict log-transformed values $\log\mu_{I}$, $\log\rho_{E}$, etc. and then apply an exponential function so that it is easier to predict parameters across a wide scale of values. We then clamp the parameters (besides $\alpha$) to the range $10^{-6},10^{6}]$; $\alpha\in(0,2)$ is enforced using a scaled sigmoid function.
Appendix H Further Details on QP Problem Classes
A summary of the problem sizes and training parameters of the different classes is presented in Table 1. For the small- to medium-scale QPs, we train all models for 500 epochs and evaluate using 1000 test samples. More training samples are used for random QPs following the setup by Saravanos et al. (2025) since the shared structure between these QPs is less clear and therefore harder to learn. As described in the main text, for the large-scale problems, we fine-tune the models trained on the smaller scale problems on 100 large-scale problems for 5 epochs, and test on 100 new problems.